I am not sure if I am supposed to subtract on the right side. I am having a hard time seeing it.
9-2m/7=19
2m/7 is a fraction.

Answer by Alan3354(69443)

9-2m/7=19
Subtract 9
-2m/7 = 10
Multiply by 7
-2m = 70
Divide by -2
m = -35